OpenStudy (anonymous):

Which of the following ratios forms a proportion with 4/15? 8/20 7/18 12/45 6/25

OpenStudy (anonymous):

8/20=2/5 7/18=7/18 12/45=4/15 6/25=6/25

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thank you